# Strong Coupling of Hydrodynamics and Reactions in Nuclear Statistical Equilibrium for Modeling Convection in Massive Stars

### v1

@@ 1 Introduction
One-dimensional models of massive star evolution leading to core-collapse can use hundreds of nuclei (e.g. Rauscher et al. 2002) including important electron/positron captures and decays (Heger et al., 2001), allowing for detailed evolution of the electron fraction in the iron core. In contrast, most multidimensional models of massive star convection use smaller networks, like the standard 21-nuclei aprox21 network, as their workhorse. These smaller networks may not be able to capture the energetics and nucleosynthesis (Navó et al., 2023).
@@
All of these works have shown that asymmetries can have important impacts on the core-collapse explosion mechanism (see also Müller 2020). Capturing the evolution of the core is difficult however, both due to the coordinate singularity in some discretizations as well as the differing timescales of the reactions in the core versus the convective shells. Additionally, many weak interactions are important in the evolution of the electron fraction in the core that cannot be captured with a simple network.
@@
In Zingale et al. (2022) (henceforth Z22), we introduced a coupling method for hydrodynamics and reactions based on the ideas of spectral deferred corrections (SDC) (Dutt et al., 2000; Minion, 2003; Zingale et al., 2019), but designed to work with the standard characteristic-tracing approach to hydrodynamics used in the corner transport upwind (CTU) (Colella, 1990) method and its extension to the piecewise parabolic method (PPM) (Colella & Woodward, 1984; Miller & Colella, 2002). We termed this integration method “simplified-SDC”. With operator splitting, when the reactions are evolved, the “pdV” expansion is not seen. This means that if a lot of heat is released into a zone in a timestep, we will not see the expansion and cooling that should accompany this. In contrast to operator splitting, in SDC the advection and reactions explicitly see the effects of each other, and in Z22, we showed that this allows us to model explosive reactive flows more accurately and efficiently.
@@ 3.1 Reaction Network and NSE
For this work, we generate the NSE table following the approach of Seitenzahl et al. (2009), as implemented in pynucastro (Smith Clark et al., 2022). Our goal with this paper is to describe the framework for coupling NSE and reactions to hydrodynamics in an accurate fashion. In a follow-on work, we will explore how the details of the nuclear physics used with the on-grid network and NSE table affect the simulations. The entire process for generating the NSE table is open, allowing for the table to easily be updated with new rates. By controlling the generation of the table ourselves, we can experiment with more recent weak rate tabulations and create custom mappings to the on-grid nuclei for a more extensive network. An overview of the table generation is given in Appendix A. Similar to other NSE tabulations, our NSE table takes as inputs $\rho$, $T$, and $Y_{e}$ and tabulates $\bar{A}$, $dY_{e}/dt$, and $\langle B/A\rangle$, $(d\langle B/A\rangle/dt)_{\mathrm{weak}}$, and $\epsilon_{\nu,\mathrm{weak}}$, where the “weak” subscript indicates that these terms are due to the weak rate evolution of the NSE state. By storing $dY_{e}/dt$, $(d\langle B/A\rangle/dt)_{\mathrm{weak}}$, and the weak rate neutrino energy loss, $\epsilon_{\nu,\mathrm{weak}}$, we are able to capture the evolution of the NSE state due to electron/positron captures and decays over a timestep. We note that $(d\bar{A}/dt)_{\mathrm{weak}}$ is very small, so we do not consider it.
In our simulations, we carry all 19 isotopes in the main network in each zone and advect them in the hydrodynamics portion of the algorithm. The composition of the larger NSE table is mapped into the 19 isotopes we carry which are also stored on the table. We’ll refer to this reduced composition as $\tilde{X}_{k}$. Because the NSE table uses many more nuclei, these approximate $\tilde{X}_{k}$ do not satisfy Eqs. 3 to 5 directly. This is why we need to explicitly carry the $\alpha_{l}$ on the grid and evolve them in the hydrodynamics system.
@@
${\tt nse\_table}(\rho,T,Y_{e})\rightarrow\bar{A},\tilde{X}_{k},\left\langle\frac{B}{A}\right\rangle,\frac{dY_{e}}{dt},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}},\epsilon_{\nu,\mathrm{weak}}$ (31)
We found that to get second-order convergence, we need to use tricubic interpolation, and use a rather dense spacing in $(\rho,T,Y_{e})$. It also helps to store $\langle B/A\rangle$ with enough precision that roundoff error does not pollute the energy generation. Ultimately, the accuracy with which nuclear masses are measured experimentally may limit the convergence.
@@ 3.2 NSE condition
For each zone, we need to be able to determine if we are in NSE. Since we integrate energy during the reactions we may begin a zone update out of NSE but enter NSE in the middle of a timestep as the temperature rises. We will address this possibility in the next section. Since we don’t have the capability to evaluate individual rates from the table, we use heuristics to determine if NSE applies. We treat a zone as being in NSE if the density and temperature exceed a threshold and the composition is mainly $\alpha$ and Fe-group nuclei (${}^{48}\mathrm{Cr}$, ${}^{52}\mathrm{Fe}$, ${}^{54}\mathrm{Fe}$, and ${}^{56}\mathrm{Ni}$). The full condition is:
@@ 4 Coupling NSE with Reaction Networks and Hydrodynamics
$\epsilon=\epsilon_{\mathrm{nuc}}-\epsilon_{\nu}-\epsilon_{\nu,\mathrm{weak}}$ (37)
Here $\epsilon_{\mathrm{nuc}}$ is the energy generation from the nuclear reactions, $\epsilon_{\nu}$ is the sum of the plasma, photo-, pair-, recombination, and Bremsstrahlung neutrino losses computed via Itoh et al. (1996), and, when using the NSE table, $\epsilon_{\nu,\mathrm{weak}}$ are the neutrino losses from electron-captures and beta-decays (and their positron equivalents). Each of these terms needs to be dealt with separately.
@@
$\frac{d(\rho e)}{dt}=-{\left[{\mathcal{{A}}}\left(\rho e\right)\right]}^{n+1/2,(k)}+\rho\epsilon_{\mathrm{nuc}}-\rho\epsilon_{\nu}$ (38)
calling the equation of state to get the temperature each time it is needed and evaluating the rates using the aprox19 network, with no contribution from $\epsilon_{\nu,\mathrm{weak}}$. We then compute $\epsilon_{\mathrm{nuc}}$ from the change in mass:
@@ 4.3 Integration Algorithm
    $\epsilon_{\nu}^{s}=\epsilon_{\nu}(\rho^{s},T^{s},\bar{A}^{s},\bar{Z}^{s})$ (51)
@@
    ${\tt nse\_table}(\rho^{s},T^{s},(Y_{e})^{s})\rightarrow\bar{A}^{s},(\tilde{X}_{k})^{s}\left\langle\frac{B}{A}\right\rangle^{s},\left[\frac{dY_{e}}{dt}\right]^{s},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{s},\epsilon_{\nu,\mathrm{weak}}^{s}$ (52)
@@
    $R^{\star}(\rho e)$ $=$ $N_{A}\rho^{s}\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{s}+N_{A}\Delta m_{np}c^{2}\rho^{s}\left[\frac{dY_{e}}{dt}\right]^{s}-\rho^{s}(\epsilon_{\nu}^{s}+\epsilon_{\nu,\mathrm{weak}}^{s})$ (53)
@@
    ${\tt nse\_table}(\rho^{\tau},T^{\tau},(Y_{e})^{\tau})\rightarrow\bar{A}^{\tau},(\tilde{X}_{k})^{\tau},\left\langle\frac{B}{A}\right\rangle^{\tau},\left[\frac{dY_{e}}{dt}\right]^{\tau},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{\tau},\epsilon_{\nu,\mathrm{weak}}^{\tau}$ (59)
@@
    $R^{s}(\rho e)$ $=$ $N_{A}\frac{\Delta(\rho\langle B/A\rangle)}{\tau}+N_{A}\Delta m_{np}c^{2}\rho^{s}\left[\frac{dY_{e}}{dt}\right]^{s}-\rho^{s}(\epsilon_{\nu}^{s}+\epsilon_{\nu,\mathrm{weak}}^{s})$ (64)
@@
This completes the estimation of the reactive sources. A typical value of $\tau$ is $\tau=\Delta t/20$.
@@
${\tt nse\_table}(\rho^{n+1},T^{n+1},(Y_{e})^{n+1})\rightarrow\bar{A}^{n+1},(\tilde{X}_{k})^{n+1},\left\langle\frac{B}{A}\right\rangle^{n+1},\left[\frac{dY_{e}}{dt}\right]^{n+1},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{n+1},\epsilon_{\nu,\mathrm{weak}}^{n+1}$ (73)
@@ 4.4 NSE Bailout
Because $\rho$ and $T$ evolve during the reactions, it is possible for a zone to start out not in NSE but evolve into NSE during the reaction update. When this happens, the ODE integrator may fail, because an excessive number of timesteps is required. To account for this during the integration, we check to see if zone evolved into NSE, and if so, we finish the zone update using the NSE prescription. To ensure that this is not triggered at the very start of integration, we require a minimum number of steps (10) to have elapsed before checking for NSE. After we leave the integrator, the state is returned to the burner driver where we finish the integration via NSE if the conditions satisfy the NSE criteria. We optionally apply a relaxation factor, typically, $0.9$, to the NSE conditions to allow a state that is close to NSE after an integrator failure to enter NSE. If a burn doesn’t enter NSE but still fails, then we trigger the Castro retry mechanism and throw away the entire timestep and start over with a smaller $\Delta t$.
@@ 5.1 NSE convergence test problem
where $r$ is the distance from the center of the domain, $L$ is the size of the domain, $r_{0}=L/2$, $u_{0}$ and $v_{0}$ are the initial $x$- and $y$-velocities. The value for each parameter is given in Table 1. We chose initial velocity values that keep the motion subsonic so the hydrodynamic slope limiters will not impact the convergence rate much. The peak Mach number initially is $\mathcal{M}=0.17$. Our initial conditions have $Y_{e}$ varying between $[0.475,0.5]$ initially. The mass fractions and remaining composition variables are then initialized from the NSE table.
@@
where $N$ is the number of zones in each direction. We run for 0.032 s with $N=32$, $64$, $128$, $256$, and $512$ zones.
Table 2 shows the convergence. We see that we get close to second order convergence for the lower resolution runs. At higher resolution, the size of the NSE table seems to be the main factor affecting the convergence. This makes sense, since, as we increase the resolution, the truncation error from the hydrodynamics will eventually get smaller than the error from interpolating in the fixed-size table. This test showed us the need for high-order interpolation and a dense NSE table. The conditions for this convergence test are representative of those that we would find in a massive star, and we note that the finest resolution used in the convergence test is $0.39~{}\mathrm{km}$, which is much finer than the typical resolution we would use with a star.
@@ 5.2.1 Initial Model
We use the same $15~{}M_{\odot}$ initial model as Fields & Couch (2020), generated by MESA (Paxton et al., 2011, 2018), representing the state of the star approximately 400 s before collapse (as determined by MESA). The MESA model is interpolated onto a uniform grid and then we reset the composition to the NSE composition if it meets the criteria Eqs. 3336. We next need to make it consistent with our equation of state and discrete form of hydrostatic equilibrium. This is done by starting at the center and using the model’s $T$ and composition (including $Y_{e}$) to find the $\rho$ and $p$ that satisfy HSE (in 1D spherical coordinates) and the EOS. For zone $i$, this has the form:
@@
Figure 1: $15M_{\odot}$ initial model showing the structure and composition of the region near the core. The dotted vertical line shows the radius inside of which we refine to higher resolution. The solid vertical line shows the radius corresponding to our domain width.

### v2

@@ 1 Introduction
One-dimensional models of massive star evolution leading to core-collapse can use hundreds of nuclei (e.g. Rauscher et al. 2002) including important electron/positron captures and decays (Heger et al., 2001), allowing for detailed evolution of the electron fraction in the iron core. In contrast, most multidimensional models of massive star convection use smaller networks, like the 21-nuclei aprox21 network, as their workhorse. These smaller networks may not be able to capture the energetics and nucleosynthesis (Farmer et al., 2016; Kato et al., 2020; Navó et al., 2023).
@@
All of these works have shown that asymmetries can have important impacts on the core-collapse explosion mechanism (see also Müller 2020). Capturing the evolution of the core is difficult however, both due to the coordinate singularity in some discretizations as well as the differing timescales of the reactions in the core versus the convective shells. Additionally, many weak interactions are important in the evolution of the electron fraction in the core that cannot be captured with a simple network (see, e.g., Farmer et al. 2016).
@@
In Zingale et al. (2022) (henceforth Z22), we introduced a coupling method for hydrodynamics and reactions based on the ideas of spectral deferred corrections (SDC) (Dutt et al., 2000; Minion, 2003; Zingale et al., 2019), but designed to work with the standard characteristic-tracing approach to hydrodynamics used in the corner transport upwind (CTU) (Colella, 1990) method and its extension to the piecewise parabolic method (PPM) (Colella & Woodward, 1984; Miller & Colella, 2002). We termed this integration method “simplified-SDC”. With operator splitting, when the reactions are evolved, the expansion of the fluid via mechanical work is not seen. This means that if a lot of heat is released into a zone in a timestep, we will not see the expansion and cooling that should accompany this. In contrast to operator splitting, in SDC the advection and reactions explicitly see the effects of each other, and in Z22, we showed that this allows us to model explosive reactive flows more accurately and efficiently.
@@ 3.1 Reaction Network and NSE
For this work, we generate the NSE table following the approach of Seitenzahl et al. (2009), as implemented in pynucastro (Smith Clark et al., 2022). Our goal with this paper is to describe the framework for coupling NSE and reactions to hydrodynamics in an accurate fashion. In a follow-on work, we will explore how the details of the nuclear physics used with the on-grid network and NSE table affect the simulations. The entire process for generating the NSE table is open source, allowing for the table to easily be updated with new rates. By controlling the generation of the table ourselves, we can experiment with more recent weak rate tabulations and create custom mappings to the on-grid nuclei for a more extensive network. An overview of the table generation is given in Appendix A. Similar to other NSE tabulations, our NSE table takes as inputs $\rho$, $T$, and $Y_{e}$ and tabulates $\bar{A}$, $dY_{e}/dt$, and $\langle B/A\rangle$, $(d\langle B/A\rangle/dt)_{\mathrm{weak}}$, and $\epsilon_{\nu,\mathrm{react}}$, where the “weak” subscript indicates that these terms are due to the weak rate evolution of the NSE state. By storing $dY_{e}/dt$, $(d\langle B/A\rangle/dt)_{\mathrm{weak}}$, and the weak rate neutrino energy loss, $\epsilon_{\nu,\mathrm{react}}$, we are able to capture the evolution of the NSE state due to electron/positron captures and decays over a timestep. We note that $(d\bar{A}/dt)_{\mathrm{weak}}$ is very small, so we do not consider it.
In our simulations, we carry all 19 isotopes in the main network in each zone and advect them in the hydrodynamics portion of the algorithm. The 96 nuclei in the NSE table are mapped into the 19 isotopes we carry which are also stored on the table. We’ll refer to this reduced composition as $\tilde{X}_{k}$. Because the NSE table uses many more nuclei, these approximate $\tilde{X}_{k}$ do not satisfy Eqs. 3 to 5 directly. This is why we need to explicitly carry the $\alpha_{l}$ on the grid and evolve them in the hydrodynamics system.
@@
${\tt nse\_table}(\rho,T,Y_{e})\rightarrow\bar{A},\tilde{X}_{k},\left\langle\frac{B}{A}\right\rangle,\frac{dY_{e}}{dt},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}},\epsilon_{\nu,\mathrm{react}}$ (31)
We found that to get second-order convergence, we need to use tricubic interpolation with each direction equivalent to a cubic Lagrange polynomial. Ultimately, the uncertainty with which nuclear masses are measured experimentally (Huang et al., 2017) may limit the convergence.
@@ 3.2 NSE condition
For each zone, we need to be able to determine if we are in NSE. Since we integrate energy during the reactions (see Eq. 24) we may begin a zone update out of NSE but enter NSE in the middle of a timestep as the temperature rises. We will address this possibility in the next section. Since we don’t have the capability to evaluate individual rates from the table, we use heuristics to determine if NSE applies. We treat a zone as being in NSE if the density and temperature exceed a threshold and the composition is mainly $\alpha$ and Fe-group nuclei (${}^{48}\mathrm{Cr}$, ${}^{52}\mathrm{Fe}$, ${}^{54}\mathrm{Fe}$, and ${}^{56}\mathrm{Ni}$). The full condition is:
@@ 4 Coupling NSE with Reaction Networks and Hydrodynamics
$\epsilon=\epsilon_{\mathrm{nuc}}-\epsilon_{\nu,\mathrm{thermal}}-\epsilon_{\nu,\mathrm{react}}$ (37)
Here $\epsilon_{\mathrm{nuc}}$ is the energy generation from the nuclear reactions, $\epsilon_{\nu,\mathrm{thermal}}$ is the sum of the plasma, photo-, pair-, recombination, and Bremsstrahlung neutrino losses computed via Itoh et al. (1996), and, when using the NSE table, $\epsilon_{\nu,\mathrm{react}}$ are the neutrino losses from electron-captures and beta-decays (and their positron equivalents). Each of these terms needs to be dealt with separately.
@@
$\frac{d(\rho e)}{dt}={\left[{\mathcal{{A}}}\left(\rho e\right)\right]}^{n+1/2,(k)}+\rho\epsilon_{\mathrm{nuc}}-\rho\epsilon_{\nu,\mathrm{thermal}}$ (38)
calling the equation of state to get the temperature each time it is needed and evaluating the rates using the aprox19 network, with no contribution from $\epsilon_{\nu,\mathrm{react}}$. We then compute $\epsilon_{\mathrm{nuc}}$ from the change in mass:
@@ 4.3 Integration Algorithm
    $\epsilon_{\nu,\mathrm{thermal}}^{s}=\epsilon_{\nu,\mathrm{thermal}}(\rho^{s},T^{s},\bar{A}^{s},\bar{Z}^{s})$ (51)
@@
    ${\tt nse\_table}(\rho^{s},T^{s},(Y_{e})^{s})\rightarrow\bar{A}^{s},(\tilde{X}_{k})^{s}\left\langle\frac{B}{A}\right\rangle^{s},\left[\frac{dY_{e}}{dt}\right]^{s},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{s},\epsilon_{\nu,\mathrm{react}}^{s}$ (52)
@@
    $R^{\star}(\rho e)$ $=$ $N_{A}\rho^{s}\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{s}+N_{A}\Delta m_{np}c^{2}\rho^{s}\left[\frac{dY_{e}}{dt}\right]^{s}-\rho^{s}(\epsilon_{\nu,\mathrm{thermal}}^{s}+\epsilon_{\nu,\mathrm{react}}^{s})$ (53)
@@
    ${\tt nse\_table}(\rho^{\tau},T^{\tau},(Y_{e})^{\tau})\rightarrow\bar{A}^{\tau},(\tilde{X}_{k})^{\tau},\left\langle\frac{B}{A}\right\rangle^{\tau},\left[\frac{dY_{e}}{dt}\right]^{\tau},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{\tau},\epsilon_{\nu,\mathrm{react}}^{\tau}$ (59)
@@
    $R^{s}(\rho e)$ $=$ $N_{A}\frac{\Delta(\rho\langle B/A\rangle)}{\tau}+N_{A}\Delta m_{np}c^{2}\rho^{s}\left[\frac{dY_{e}}{dt}\right]^{s}-\rho^{s}(\epsilon_{\nu,\mathrm{thermal}}^{s}+\epsilon_{\nu,\mathrm{react}}^{s})$ (64)
@@
This completes the estimation of the reactive sources. A typical value of $\tau$ is $\tau=\Delta t/20$, which is large enough for advection to cause a change in the thermodynamic state, but overall the convergence result appears insensitive to this choice. This is a runtime option in the code.
@@
${\tt nse\_table}(\rho^{n+1},T^{n+1},(Y_{e})^{n+1})\rightarrow\bar{A}^{n+1},(\tilde{X}_{k})^{n+1},\left\langle\frac{B}{A}\right\rangle^{n+1},\left[\frac{dY_{e}}{dt}\right]^{n+1},\left[\frac{d\langle B/A\rangle}{dt}\right]_{\mathrm{weak}}^{n+1},\epsilon_{\nu,\mathrm{react}}^{n+1}$ (73)
@@ 4.4 NSE Bailout
Because $\rho$ and $T$ evolve during the reactions, it is possible for a zone to start out not in NSE but evolve into NSE during the reaction update. When this happens, the ODE integrator may fail, because an excessive number of timesteps is required. To account for this during the integration, we check to see if zone evolved into NSE, and if so, we finish the zone update using the NSE prescription. To ensure that this is not triggered at the very start of integration, we require a minimum number of steps (10) to have elapsed before checking for NSE (this is enough to get past the timestep estimation stage in VODE that occupies the first timesteps). After we leave the integrator, the state is returned to the burner driver where we finish the integration via NSE if the conditions satisfy the NSE criteria. We optionally apply a relaxation factor (emperically suggested to be $0.9$) to the NSE conditions to allow a state that is close to NSE after an integrator failure to enter NSE. If a burn doesn’t enter NSE but still fails, then we trigger the Castro retry mechanism and throw away the entire timestep and start over with a smaller $\Delta t$.
@@ 5.1 NSE convergence test problem
where $r$ is the distance from the center of the domain, $L$ is the size of the domain, $r_{0}=L/2$, $u_{0}$ and $v_{0}$ are the initial $x$- and $y$-velocities. The value for each parameter is given in Table 1. We chose initial velocity values that keep the motion subsonic, minimizing pressure fluctuations, so the hydrodynamic slope limiters will not reduce the order of the parabolic reconstruction and impact the convergence rate. The peak Mach number initially is $\mathcal{M}=0.17$. Our initial conditions have $Y_{e}$ varying between $[0.475,0.5]$ initially. The mass fractions and remaining composition variables are then initialized from the NSE table.
@@
where $N$ is the number of zones in each direction. We run for 0.032 s with $N=32$, $64$, $128$, $256$, and $512$ zones. Since the timestep is reduced in direct proportion to the grid spacing, this convergence test simultaneously tests convergence in space and time.
For a pair of runs with $N$ and $2N$ zones, we coarsen the finer resolution simulation via conservative averaging down to $N$ zones (we call this $\phi^{(2N,\mathrm{coarsened})}$) and compute the $L_{1}$ norm of the difference with the simulation run with $N$ zones, $\phi^{(N)}$ as:
$\epsilon_{N\rightarrow 2N}=\|\phi^{(N)}-\phi^{(2N,\mathrm{coarsened})}\|_{1}=\Delta x\Delta y\sum_{i,j}|\phi^{(N)}_{i,j}-\phi^{(2N,\mathrm{coarsened})}_{i,j}|$ (82)
We can then compute the rate of convergence, $r$, as:
$r=\log_{2}\frac{\epsilon_{N\rightarrow 2N}}{\epsilon_{2N\rightarrow 4N}}$ (83)
Table 2 shows the convergence. For the 5 different resolutions, we can measure 3 convergence rates. We do this for several state variables.
We see that we get close to second order convergence for the lower resolution runs. At higher resolution, the size of the NSE table seems to be the main factor affecting the convergence—with a more widely spaced tabulation than used here, we observed worse convergence. This makes sense, since, as we increase the resolution, the truncation error from the hydrodynamics will eventually get smaller than the error from interpolating in the fixed-size table. This test showed us the need for high-order interpolation and a dense NSE table. The conditions for this convergence test are representative of those that we would find in a massive star, and we note that the finest resolution used in the convergence test is $0.39~{}\mathrm{km}$, which is much finer than the typical resolution we would use with a star, as we show in section 5.2.
@@ 5.2.1 Initial Model
We use the same $15~{}M_{\odot}$ initial model as Fields & Couch (2020), generated by MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019), representing the state of the star approximately 400 s before collapse (as determined by MESA). At this point in time, the star’s mass is about $13.3~{}M_{\odot}$. The MESA model is interpolated onto a uniform grid and then we reset the composition to the NSE composition if it meets the criteria Eqs. 3336. We next need to make it consistent with our equation of state and discrete form of hydrostatic equilibrium. This is done by starting at the center and using the model’s $T$ and composition (including $Y_{e}$) to find the $\rho$ and $p$ that satisfy HSE (in 1D spherical coordinates) and the EOS. For zone $i$, this has the form:
@@
Figure 1: $15M_{\odot}$ initial model (from Fields & Couch (2020)) showing the structure and composition of the region near the core. The dotted vertical line shows the radius inside of which we refine to higher resolution. The solid vertical line shows the radius corresponding to our domain width.